# Orbifold modifications of complex analytic varieties

### v4

@@ 4.C. Equations of standard form
$\varphi\colon(x,y,z,s)\mapsto(a,b,c,s^{\prime}),$
@@
[…]
@@
Let $\tau\colon X^{\prime}\to X$ be the projective bimeromorphic morphism induced by the functorial desingularization of $S$, as shown at the beginning of Subsection 3.B. Then, by replacing $X$ with $X^{\prime}$, $S$ with its strict transform $S^{\prime}$ in $X^{\prime}$, and $C$ with the union of its preimage in $S^{\prime}$ and the intersection of $S^{\prime}$ with the $\tau$-exceptional locus, we may assume that $S$ is smooth. By applying Corollary 4.6 and then Lemma 4.7, we may assume that $X$ has double-point singularities along $S$.
@@
We only treat the case when $F_{2}\neq 0$. The other case can be proved by the same method. Let $p\colon X^{\prime}\to X$ be the projective bimeromorphic morphism induced by the blowup of $S$ at $\mathcal{I}$. Let $S^{\prime}$ be the strict transform of $S$ in $X^{\prime}$, and let $E$ be the divisor in $S^{\prime}$ defined by $(p|_{S^{\prime}})^{*}\mathcal{I}$. Then, as shown in Subsection 3.B, for any point $o^{\prime}$ in $S^{\prime}$, there is a neighborhood $U^{\prime}\subseteq X^{\prime}$ of $o^{\prime}$, such that $U^{\prime}$ is the hypersurface in $\mathbb{D}^{3}\times(S^{\prime}\cap U^{\prime})$, defined by the function
@@
We denote by $\mathcal{I}^{\prime}$ the ideal sheaf of the closed subset $E\cup(p|_{S^{\prime}})^{-1}(C)$ in $S^{\prime}$. Let $Y\to X^{\prime}$ be the projective bimeromorphic morphism induced by the functorial principalization of $\mathcal{I}^{\prime}$ on $S^{\prime}$, and let $f\colon Y\to X$ be the natural morphism. We denote by $S_{Y}$ the strict transform of $S^{\prime}$ in $Y$, and by $H_{Y}\subseteq S_{Y}$ the preimage of $E\cup(p|_{S^{\prime}})^{-1}(C)$. Then $H_{Y}$ can be seen as a reduced divisor in $S_{Y}$ and $(S_{Y},H_{Y})$ is a snc pair. Furthermore, for any point $o\in S_{Y}$, there is a neighborhood $U\subseteq Y$ of $o$, such that $U$ is a hypersurface in $\mathbb{D}^{3}\times(S_{Y}\cap U)$ defined by an equation of standard form with respect to $(S_{Y}\cap U,H_{Y}\cap U)$. In addition, $Y$ has the same type of canonical singularities at every point of $S_{Y}\setminus H_{Y}$, and $H_{Y}$ contains the intersection of $S_{Y}$ with the $f$-exceptional locus. This completes the proof of the proposition. ∎
@@ 5. Further improvement of defining equations
    Assume that $X$ has $A$-type or $E$-type singularities at points of $S\setminus H$. Let $f\colon Y\to X$ be the composition of a sequence of blowups, such that each center is a component $H_{i}$ of $H$, see Subsection 3.C. In particular, we can identify $S$ with its strict transform in $Y$. Assume that for each $i=1,...,n$, there are at least $N$ blowups in the sequence whose centers are $H_{i}$. Then there is a neighborhood $U\subseteq Y$ of $S\subseteq Y$, such that $U$ is isomorphic to a neighborhood of $\{\mathbf{0}_{3}\}\times S$ in the hypersurface in $\mathbb{D}^{3}\times S$ defined by a function of one of the following shapes
@@
[…]
@@ 5.A. Factorization lemmas
Then $-1+\delta(y,z,s)$ is divisible by $p(y,z,s)$. We set
@@
$H(y,z,s)$ ::=:= $(y+C(s)\cdot z^{2})(y^{2}+z^{3}-C(s)\cdot z^{2}y+C(s)^{2}\cdot z^{4})$
[…]
@@
where $a(u,s)=-3C(s)u$, and $n(u,s)=(1+3C(s)^{2}u)$ is a unit. Hence we can apply Lemma 5.3 to $\overline{F}(y^{\prime},z,s,u)$, and deduce a factorization $\overline{F}=\overline{G}_{1}\overline{G}_{2}$ with $\overline{G}_{1}={y^{\prime}}+\overline{c}({y^{\prime}},z,s)$ such that $\overline{c}$ is divisible by $z^{2}$. Furthermore, we have
@@ 5.B. Case of $A$-type singularities
By assumption, there is some integer $l\geq 1$ such that $X$ has $A_{l}$-type singularities at any point $s\in S\setminus H$. Hence $e=l+1$, and if $b(s)$ is the term in $R_{3}$ of degree 00 in $z$, then the zero locus of $b(s)$ is contained in $H$. Thus, by blowing up $X$ at the $H_{i}$’s, we may assume that $b(s)$ divides $R_{3}(z,s)$, see Remark 3.7. Then the defining equation is of the shape
@@ 5.C. Case of $E$-type singularities
such that either $v$ is zero or $z$ does not divide $v$. Moreover, the terms of $v(z,s)$ have degrees at most $q-p$ in $z$. In addition, either $w(y,z,s)$ is zero, or $w(0,0,s)\neq 0$ for a general point $s\in S$.
@@
where $q\geq p\geq 0$ are integers. Either $v$ is zero or $z$ does not divide $v$. Furthermore, the terms of $v(z,s)$ have degrees at most $q-p$ in $z$. In addition, either $w(y,z,s)$ is zero, or $w(0,0,s)\neq 0$ for a general point $s\in S$.
@@
where $q\geq p\geq 0$ are integers. Either $v$ is zero or $z$ does not divide $v$. Furthermore, the terms of $v(z,s)$ have degrees at most $q-p$ in $z$. In addition, either $w(y,z,s)$ is zero, or $w(0,0,s)\neq 0$ for a general point $s\in S$.
@@
Since $q\geq 1$, the RHS above tends to +$\infty$+ ∞ if $e_{i}$ tends to +$\infty$+ ∞. Hence we can pick $e_{i}$ sufficiently large so that
@@ 5.D. Case of $D$-type singularities
[…]
@@
Let $R_{9}(y,z)$ be a $(l-2)$-root of $R_{8}(y,s)$. Then we have
@@
such that the zero locus of $q$ and $r$ are contained in $H$, and that the zero locus of $q(s)-r(s)$ is also contained in $H$. Hence, up to switching $q$ and $r$, there are unit functions $\overline{q}$ and $\overline{r}$ such that.
@@
[…]
@@
The objective that we first construct $f_{1}\colon X_{1}\to X$ is to make the finite cover $f_{2}\colon X_{2}\to X_{1}$ functorial. For the holomorphic function $T_{i}$ on $S$, there is no canonical way to make it a holomorphic function on $X$. And the Cartier divisors in $X$ defined by $T_{i}=0$ depend on the choice of coordinates in Setup 5.1. However, in a neighborhood $V$ of the strict transform $(f_{1}^{-1})_{*}S$ of $S$ in $X_{1}$, the Cartier divisor defined by $T_{i}=0$ is the $f_{1}$-exceptional divisor defined by $f_{1}^{-1}\mathcal{I}({H_{i}})\cdot\mathcal{O}_{X_{1}}$. In particular, this Cartier divisor is intrinsic. To see this more explicitly, we assume that
$\varphi\colon(x,y,z,s)\mapsto(x^{\prime},y^{\prime},z^{\prime},s^{\prime})$
@@
$f_{1}\colon(x_{1}^{\prime},y_{1}^{\prime},z_{1}^{\prime},s^{\prime})\mapsto(T(s^{\prime})x_{1}^{\prime},T(s^{\prime})y_{1}^{\prime},T(s^{\prime})z_{1}^{\prime},s^{\prime})=(x^{\prime},y^{\prime},z^{\prime},s^{\prime})$
As a consequence, the change of coordinates $\varphi_{1}\colon(x_{1},y_{1},z_{1},s)\mapsto(x_{1}^{\prime},y_{1}^{\prime},z_{1}^{\prime},s^{\prime})$ satisfies
@@
As a result, by Lemma 2.5, the finite morphism $f_{2}\colon X_{2}\to X_{1}$ constructed via two choices of coordinates in Setup 5.1 are the same on $V$.
@@ 6. Local construction of covering spaces
If $\overline{m}\in M$ is another point, then there are isomorphisms $\pi_{1}(M,m)\cong\pi_{1}(M,\overline{m})$ induced by the paths joining $m$ and $\overline{m}$. These isomorphisms are in bijection with the inner automorphisms of $\pi_{1}(M,\overline{m}).$ Now we assume that $f\colon M\to M^{\prime}$ is a continuous map between path-connected topological spaces. If $m^{\prime}=f(m)$, then there is a natural morphism
$f_{m,*}\colon\pi_{1}(M,m)\to\pi_{1}(M^{\prime},m^{\prime}),$
which induces a morphism of fundamental groups $\pi_{1}(M)\to\pi_{1}(M^{\prime})$. The conjugacy class of $\pi_{1}(M)\to\pi_{1}(M^{\prime})$ is independent of the choice of basepoints. By abuse of notation, we will call this conjugacy class the natural morphism induced by $f$, and denote it by
$f_{*}\colon\pi_{1}(M)\to\pi_{1}(M^{\prime}).$
@@ 6.A. Preparatory lemmas
Let $U\subseteq\mathbb{C}^{m}$ and $S\subseteq\mathbb{C}^{n}$ be open neighborhoods of the origins, and let $f\colon U\times S\to S$ be a morphism such that $f(\mathbf{0}_{m},\mathbf{0}_{n})=\mathbf{0}_{n}.$ Then there are open neighborhoods $W\subseteq U$ and $T\subseteq S$ of the origins, such that $f|_{W\times T}$ is a continuous deformation equivalence to the projection $p$ onto $T$, with values contained in $S$. In other words, there is a continuous function
$H(u,s,\mu)\colon W\times T\times[0,1]\to S,$
@@
There is a real number $r>0$ such that $S$ contains the ball in $\mathbb{C}^{n}$ of radius $02r>$ 0 centered at the origin. We can define the function $g(u,s)=f(u,s)-s$ on $U\times S$, which takes values in $\mathbb{C}^{n}$. Then the assumption implies that $g(\mathbf{0}_{m},\mathbf{0}_{n})=\mathbf{0}_{n}.$ Thus there are open neighborhoods $W\subseteq U$ and $T\subseteq S$ of the origins, such that $|g(u,s)|<r$ for $(u,s)\in W\times T$. We can choose $T$ so that it is contained in the ball in $\mathbb{C}^{n}$ of radius $r$ centered at the origin. We define the function $H(u,s,\mu)=s+\mu g(u,s)$ on $W\times T\times[0,1]$. Then $H$ has values in $S$, and it induces a deformation equivalence from $f|_{W\times T}$ to $p$. This completes the proof of the lemma. ∎
@@
$(\varphi_{0})_{*},(\varphi_{1})_{*}\colon\pi_{1}(M)\to\pi_{1}(M^{\prime})$
@@
For $i=0,1$, we let $\iota_{i}\colon M\to M\times[0,1]$ be the continuous map sending a point $m\in M$ to $(m,i)$. Then both $\iota_{0}$ and $\iota_{1}$ induce an isomorphism from $\pi_{1}(M)$ to $\pi_{1}(M\times I)$. Since $\varphi_{i}=H\circ\iota_{i}$ for $i=0,1$, the images of $(\varphi_{0})_{*}$ and $(\varphi_{1})_{*}$ are both equal to the image of $H_{*}\colon\pi_{1}(M\times I)\to\pi_{1}(M^{\prime})$. ∎
@@
$\varphi\colon(x,y,z,s)\mapsto(T^{p}x,T^{q}y,T^{r}z,s)$
@@
There three unknowns and two equations, hence the system is indeterminate and admits infinitely many integer solutions. Furthermore, since $l\geq 2$, we see that the solutions for $q,r$ tend to +$\infty$+ ∞ if $p$ tends to +$\infty$+ ∞. Hence there are integers $p,q,r\geq N_{0}$ such that $\varphi$ is well-defined. Since $T$ is nowhere vanishing on $\check{S}^{\circ}$, we deduce that $\varphi$ is an automorphism of $Y\times\check{S}^{\circ}$ over $\check{S}^{\circ}$. ∎
@@
With the notation in Lemma 6.3, Let $U\subseteq Y$ be an open neighborhood of $\mathbf{0}_{3}$ such that the natural morphism of fundamental groups $\pi_{1}(U_{\operatorname{\textsubscript{\rm sm}}})\to\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}})$ is an isomorphism. Let $S\subseteq\check{S}$ be an open subset, such that $S$ is also isomorphic to a contractible open neighborhood of the origin in $\mathbb{C}^{n}$, and that $H|_{S}$ is isomorphic to the union of coordinates hyperplanes. Let $S^{\circ}=S\setminus H$.
Let $\varphi\colon U\times S^{\circ}\to Y\times\check{S}^{\circ}$ be an open embedding, whose restriction on $\{\mathbf{0}_{3}\}\times S^{\circ}$ is the identity map. Then the induced morphism $\varphi_{*}$ of fundamental groups from $\pi_{1}(U_{\operatorname{\textsubscript{\rm sm}}}\times S^{\circ})$ to $\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$ is injective. Furthermore, its image is equal to the image of $\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times S^{\circ})$ induced by the natural inclusion. In particular, if $k^{\prime}$ is the number of irreducible components of $H|_{S}$, then we have
@@
$\varphi\colon(x,y,z,s)\mapsto(x^{\prime},y^{\prime},z^{\prime},s^{\prime})$
@@
$\psi\colon(x,y,z,s)\mapsto(x,y,z,s^{\prime})$
@@
By Lemma 6.1, there is an open neighborhood $W$ of $\mathbf{0}_{3}$ in $\overline{W}$, there are open neighborhoods $T\subseteq T^{\prime}$ of $o$ in $\overline{T}$, such that the restriction of $s^{\prime}$ on $W\times T$ is deformation equivalent to the projection function $(x,y,z,s)\mapsto s$, with values inside $T^{\prime}$. Up to shrinking $W$ and $T^{\prime}$, we may assume that $W\times T^{\prime}\subseteq Z$. As a consequence, the image $\psi(W_{sm}\times\{o\})$ is deformation equivalent to $W_{\operatorname{\textsubscript{\rm sm}}}\times\{o\}$ inside $Z_{\operatorname{\textsubscript{\rm sm}}}=Z\setminus(\{\mathbf{0}_{3}\}\times S^{\circ})$. Then by Lemma 6.2, the image of
$\psi_{*}\colon\pi_{1}(W_{\operatorname{\textsubscript{\rm sm}}}\times\{o\})\to\pi_{1}(Z_{\operatorname{\textsubscript{\rm sm}}})$

### v6

@@ 4.C. Equations of standard form
$\varphi\colon(x,y,z,s)\mapstochar\rightarrow(a,b,c,s^{\prime}),$
@@
Comparing the previous equation with (4.1) and (4.4), we deduce that
@@
Let $\tau\colon X^{\prime}\rightarrow X$ be the projective bimeromorphic morphism induced by the functorial desingularization of $S$, as shown at the beginning of Subsection 3.B. Then, by replacing $X$ with $X^{\prime}$, $S$ with its strict transform $S^{\prime}$ in $X^{\prime}$, and $C$ with the union of its preimage in $S^{\prime}$ and the intersection of $S^{\prime}$ with the $\tau$-exceptional locus, we may assume that $S$ is smooth. By applying Corollary 4.6 and then Lemma 4.7, we may assume that $X$ has double-point singularities along $S$.
@@
We only treat the case when $F_{2}\neq 0$. The other case can be proved by the same method. Let $p\colon X^{\prime}\rightarrow X$ be the projective bimeromorphic morphism induced by the blowup of $S$ at $\mathcal{I}$. Let $S^{\prime}$ be the strict transform of $S$ in $X^{\prime}$, and let $E$ be the divisor in $S^{\prime}$ defined by $(p|_{S^{\prime}})^{*}\mathcal{I}$. Then, as shown in Subsection 3.B, for any point $o^{\prime}$ in $S^{\prime}$, there is a neighborhood $U^{\prime}\subseteq X^{\prime}$ of $o^{\prime}$, such that $U^{\prime}$ is the hypersurface in $\mathbb{D}^{3}\times(S^{\prime}\cap U^{\prime})$, defined by the function
@@
We denote by $\mathcal{I}^{\prime}$ the ideal sheaf of the closed subset $E\cup(p|_{S^{\prime}})^{-1}(C)$ in $S^{\prime}$. Let $Y\rightarrow X^{\prime}$ be the projective bimeromorphic morphism induced by the functorial principalization of $\mathcal{I}^{\prime}$ on $S^{\prime}$, and let $f\colon Y\rightarrow X$ be the natural morphism. We denote by $S_{Y}$ the strict transform of $S^{\prime}$ in $Y$, and by $H_{Y}\subseteq S_{Y}$ the preimage of $E\cup(p|_{S^{\prime}})^{-1}(C)$. Then $H_{Y}$ can be seen as a reduced divisor in $S_{Y}$ and $(S_{Y},H_{Y})$ is a snc pair. Furthermore, for any point $o\in S_{Y}$, there is a neighborhood $U\subseteq Y$ of $o$, such that $U$ is a hypersurface in $\mathbb{D}^{3}\times(S_{Y}\cap U)$ defined by an equation of standard form with respect to $(S_{Y}\cap U,H_{Y}\cap U)$. In addition, $Y$ has the same type of canonical singularities at every point of $S_{Y}\setminus H_{Y}$, and $H_{Y}$ contains the intersection of $S_{Y}$ with the $f$-exceptional locus. This completes the proof of the proposition. ∎
@@ 5. Further improvement of defining equations
    Assume that $X$ has $A$-type or $E$-type singularities at points of $S\setminus H$. Let $f\colon Y\rightarrow X$ be the composition of a sequence of blowups, such that each center is a component $H_{i}$ of $H$, see Subsection 3.C. In particular, we can identify $S$ with its strict transform in $Y$. Assume that for each i=1,,ni=1,...,n, there are at least $N$ blowups in the sequence whose centers are $H_{i}$. Then there is a neighborhood $U\subseteq Y$ of $S\subseteq Y$, such that $U$ is isomorphic to a neighborhood of $\{\mathbf{0}_{3}\}\times S$ in the hypersurface in $\mathbb{D}^{3}\times S$ defined by a function of one of the following shapes
@@
    Assume that $X$ has $D$-type singularities at points of $S\setminus H$. Let $f_{1}\colon X_{1}\rightarrow X$ be the composition of blowups at the $H_{i}$’s as in Subsection 3.C, such that for each i=1,,ni=1,...,n, there is exactly one blowup whose center is $H_{i}$. Let $f_{2}\colon X_{2}\rightarrow X_{1}$ be the basechange over $S[\sqrt[6]{T_{1}},...,\sqrt[6]{T_{n}}]$. We define $\overline{S}=f_{2}^{-1}((f_{1})^{-1}_{*}S)$ and define $\overline{H}_{i}$ as the preimage of $H_{i}$ in $\overline{S}$. Let $f_{3}\colon Y\rightarrow X_{2}$ be the composition of a sequence of blowups at the $\overline{H}_{i}$’s as in Subsection 3.C. Assume that for each i=1,,ni=1,...,n, there are at least $N$ blowups in the sequence whose centers are $\overline{H}_{i}$. Let $f\colon Y\rightarrow X$ be the natural morphism, let $\overline{S}_{Y}$ be the strict transform of $\overline{S}$ in $Y$. Then there is a neighborhood $U\subseteq Y$ of $\overline{S}_{Y}$, such that $U$ is isomorphic to a neighborhood of $\{\mathbf{0}_{3}\}\times\overline{S}_{Y}$ in the hypersurface in $\mathbb{D}^{3}\times\overline{S}_{Y}$ defined by a function of the shape
@@ 5.A. Factorization lemmas
Then $1+\delta(y,z,s)$ is divisible by $p(y,z,s)$. We set
@@
$H(y,z,s)$ $:=$ $(y+C(s)\cdot z^{2})(y^{2}+z^{3}-C(s)\cdot z^{2}y+C(s)^{2}\cdot z^{4})$
$\ =$ $y(y^{2}+z^{3})+C(s)\cdot z^{5}+C(s)^{3}\cdot z^{6}$
$\ =$ $F(y,z,s)+z^{5}\cdot(C(s)^{3}z-R_{1}(y,z,s))$
@@
where a(u,s)=3C(s)ua(u,s)=-3C(s)u, and $n(u,s)=(1+3C(s)^{2}u)$ is a unit. Hence we can apply Lemma 5.3 to $\overline{F}(y^{\prime},z,s,u)$, and deduce a factorization $\overline{F}=\overline{G}_{1}\overline{G}_{2}$ with $\overline{G}_{1}={y^{\prime}}+\overline{c}({y^{\prime}},z,s)$ such that $\overline{c}$ is divisible by $z^{2}$. Furthermore, we have
@@ 5.B. Case of $A$-type singularities
By assumption, there is some integer $l\geq 1$ such that $X$ has $A_{l}$-type singularities at any point $s\in S\setminus H$. Hence $e=l+1$, and if $b(s)$ is the term in $R_{3}$ of degree 0 in $z$, then the zero locus of $b(s)$ is contained in $H$. Thus, by blowing up $X$ at the $H_{i}$’s, we may assume that $b(s)$ divides $R_{3}(z,s)$, see Remark 3.7. Then the defining equation is of the shape
@@ 5.C. Case of $E$-type singularities
such that either $v$ is zero or $z$ does not divide $v$. Moreover, the terms of $v(z,s)$ have degrees at most qpq-p in $z$. In addition, either $w(y,z,s)$ is zero, or $w(0,0,s)\neq 0$ for a general point $s\in S$.
@@
where $q\geq p\geq 0$ are integers. Either $v$ is zero or $z$ does not divide $v$. Furthermore, the terms of $v(z,s)$ have degrees at most qpq-p in $z$. In addition, either $w(y,z,s)$ is zero, or $w(0,0,s)\neq 0$ for a general point $s\in S$.
@@
where $q\geq p\geq 0$ are integers. Either $v$ is zero or $z$ does not divide $v$. Furthermore, the terms of $v(z,s)$ have degrees at most qpq-p in $z$. In addition, either $w(y,z,s)$ is zero, or $w(0,0,s)\neq 0$ for a general point $s\in S$.
@@
Since $q\geq 1$, the RHS above tends to $\infty$ if $e_{i}$ tends to $\infty$. Hence we can pick $e_{i}$ sufficiently large so that
@@ 5.D. Case of $D$-type singularities
Let $X_{2}\rightarrow X$ be the basechange over $\overline{S}=S[\sqrt[6]{T_{1}},...,\sqrt[6]{T_{n}}]\rightarrow S$, let $\overline{S}$ be the preimage of $S$ in $X_{2}$, and let $\overline{H}_{i}$ be the preimage of $H_{i}$ for every i=1,,ni=1,...,n. In the remainder of the proof, by abuse of notation, we will write $s\in S$ for $\overline{s}\in\overline{S}$, $H=\sum_{i=1}^{n}H_{i}$ for $\overline{H}=\sum_{i=1}^{n}\overline{H}_{i}$, $X$ for a neighborhood of $\overline{S}$ in $X_{2}$, and $F$ for a defining function of $X$. Then Lemma 2.9 implies that
@@
Let $R_{9}(y,z)$ be a (l2)(l-2)-root of $R_{8}(y,s)$. Then we have
@@
such that the zero locus of $q$ and $r$ are contained in $H$, and that the zero locus of q(s)r(s)q(s)-r(s) is also contained in $H$. Hence, up to switching $q$ and $r$, there are unit functions $\overline{q}$ and $\overline{r}$ such that.
@@
If we let $\theta\colon X\rightarrow S$ be the projection induced by the canonical one from $\mathbb{D}^{3}\times S$ to $S$, then we can regard $T_{i}$ as a holomorphic function on $X_{1}$ by pulling it back via $\theta\circ f_{1}\colon X_{1}\rightarrow S$. Then in a neighborhood of the strict transform $(f_{1}^{-1})_{*}S$ of $S$ in $X_{1}$, the Cartier divisor defined by $T_{i}=0$ is the $f_{1}$-exceptional divisor defined by $f_{1}^{-1}\mathcal{I}({H_{i}})\cdot\mathcal{O}_{X_{1}}$, where $\mathcal{I}({H_{i}})$ is the ideal sheaf of $H_{i}$ in $X$, see Remark 3.5. Furthermore, we have $X_{2}=X_{1}[\sqrt[6]{T_{1}},...,\sqrt[6]{T_{n}}]$.
@@
The objective that we first construct $f_{1}\colon X_{1}\rightarrow X$ is to make the finite cover $f_{2}\colon X_{2}\rightarrow X_{1}$ functorial. For the holomorphic function $T_{i}$ on $S$, there is no canonical way to make it a holomorphic function on $X$. And the Cartier divisors in $X$ defined by $T_{i}=0$ depend on the choice of coordinates in Setup 5.1. However, in a neighborhood $V$ of the strict transform $(f_{1}^{-1})_{*}S$ of $S$ in $X_{1}$, the Cartier divisor defined by $T_{i}=0$ is the $f_{1}$-exceptional divisor defined by $f_{1}^{-1}\mathcal{I}({H_{i}})\cdot\mathcal{O}_{X_{1}}$. In particular, this Cartier divisor is intrinsic. To see this more explicitly, we assume that
$\varphi\colon(x,y,z,s)\mapstochar\rightarrow(x^{\prime},y^{\prime},z^{\prime},s^{\prime})$
@@
$f_{1}\colon(x_{1}^{\prime},y_{1}^{\prime},z_{1}^{\prime},s^{\prime})\mapstochar\rightarrow(T(s^{\prime})x_{1}^{\prime},T(s^{\prime})y_{1}^{\prime},T(s^{\prime})z_{1}^{\prime},s^{\prime})=(x^{\prime},y^{\prime},z^{\prime},s^{\prime})$
As a consequence, the change of coordinates $\varphi_{1}\colon(x_{1},y_{1},z_{1},s)\mapstochar\rightarrow(x_{1}^{\prime},y_{1}^{\prime},z_{1}^{\prime},s^{\prime})$ satisfies
@@
As a result, by Lemma 2.5, the finite morphism $f_{2}\colon X_{2}\rightarrow X_{1}$ constructed via two choices of coordinates in Setup 5.1 are the same on $V$.
@@ 6. Local construction of covering spaces
If $\overline{m}\in M$ is another point, then there are isomorphisms $\pi_{1}(M,m)\cong\pi_{1}(M,\overline{m})$ induced by the paths joining $m$ and $\overline{m}$. These isomorphisms are in bijection with the inner automorphisms of $\pi_{1}(M,\overline{m}).$ Now we assume that $f\colon M\rightarrow M^{\prime}$ is a continuous map between path-connected topological spaces. If $m^{\prime}=f(m)$, then there is a natural morphism
$f_{m,*}\colon\pi_{1}(M,m)\rightarrow\pi_{1}(M^{\prime},m^{\prime}),$
which induces a morphism of fundamental groups $\pi_{1}(M)\rightarrow\pi_{1}(M^{\prime})$. The conjugacy class of $\pi_{1}(M)\rightarrow\pi_{1}(M^{\prime})$ is independent of the choice of basepoints. By abuse of notation, we will call this conjugacy class the natural morphism induced by $f$, and denote it by
$f_{*}\colon\pi_{1}(M)\rightarrow\pi_{1}(M^{\prime}).$
@@ 6.A. Preparatory lemmas
Let $U\subseteq\mathbb{C}^{m}$ and $S\subseteq\mathbb{C}^{n}$ be open neighborhoods of the origins, and let $f\colon U\times S\rightarrow S$ be a morphism such that $f(\mathbf{0}_{m},\mathbf{0}_{n})=\mathbf{0}_{n}.$ Then there are open neighborhoods $W\subseteq U$ and $T\subseteq S$ of the origins, such that $f|_{W\times T}$ is a continuous deformation equivalence to the projection $p$ onto $T$, with values contained in $S$. In other words, there is a continuous function
$H(u,s,\mu)\colon W\times T\times[0,1]\rightarrow S,$
@@
There is a real number $r>0$ such that $S$ contains the ball in $\mathbb{C}^{n}$ of radius $2r>0$ centered at the origin. We can define the function g(u,s)=f(u,s)sg(u,s)=f(u,s)-s on $U\times S$, which takes values in $\mathbb{C}^{n}$. Then the assumption implies that $g(\mathbf{0}_{m},\mathbf{0}_{n})=\mathbf{0}_{n}.$ Thus there are open neighborhoods $W\subseteq U$ and $T\subseteq S$ of the origins, such that $|g(u,s)|<r$ for $(u,s)\in W\times T$. We can choose $T$ so that it is contained in the ball in $\mathbb{C}^{n}$ of radius $r$ centered at the origin. We define the function $H(u,s,\mu)=s+\mu g(u,s)$ on $W\times T\times[0,1]$. Then $H$ has values in $S$, and it induces a deformation equivalence from $f|_{W\times T}$ to $p$. This completes the proof of the lemma. ∎
@@
$(\varphi_{0})_{*},(\varphi_{1})_{*}\colon\pi_{1}(M)\rightarrow\pi_{1}(M^{\prime})$
@@
For $i=0,1$, we let $\iota_{i}\colon M\rightarrow M\times[0,1]$ be the continuous map sending a point $m\in M$ to $(m,i)$. Then both $\iota_{0}$ and $\iota_{1}$ induce an isomorphism from $\pi_{1}(M)$ to $\pi_{1}(M\times I)$. Since $\varphi_{i}=H\circ\iota_{i}$ for $i=0,1$, the images of $(\varphi_{0})_{*}$ and $(\varphi_{1})_{*}$ are both equal to the image of $H_{*}\colon\pi_{1}(M\times I)\rightarrow\pi_{1}(M^{\prime})$. ∎
@@
$\varphi\colon(x,y,z,s)\mapstochar\rightarrow(T^{p}x,T^{q}y,T^{r}z,s)$
@@
There three unknowns and two equations, hence the system is indeterminate and admits infinitely many integer solutions. Furthermore, since $l\geq 2$, we see that the solutions for $q,r$ tend to $\infty$ if $p$ tends to $\infty$. Hence there are integers $p,q,r\geq N_{0}$ such that $\varphi$ is well-defined. Since $T$ is nowhere vanishing on $\check{S}^{\circ}$, we deduce that $\varphi$ is an automorphism of $Y\times\check{S}^{\circ}$ over $\check{S}^{\circ}$. ∎
@@
With the notation in Lemma 6.3, Let $U\subseteq Y$ be an open neighborhood of $\mathbf{0}_{3}$ such that the natural morphism of fundamental groups $\pi_{1}(U_{\operatorname{\textsubscript{\rm sm}}})\rightarrow\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}})$ is an isomorphism. Let $S\subseteq\check{S}$ be an open subset, such that $S$ is also isomorphic to a contractible open neighborhood of the origin in $\mathbb{C}^{n}$, and that $H|_{S}$ is isomorphic to the union of coordinates hyperplanes. Let $S^{\circ}=S\setminus H$.
Let $\varphi\colon U\times S^{\circ}\rightarrow Y\times\check{S}^{\circ}$ be an open embedding, whose restriction on $\{\mathbf{0}_{3}\}\times S^{\circ}$ is the identity map. Then the induced morphism $\varphi_{*}$ of fundamental groups from $\pi_{1}(U_{\operatorname{\textsubscript{\rm sm}}}\times S^{\circ})$ to $\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times\check{S}^{\circ})$ is injective. Furthermore, its image is equal to the image of $\pi_{1}(Y_{\operatorname{\textsubscript{\rm sm}}}\times S^{\circ})$ induced by the natural inclusion. In particular, if $k^{\prime}$ is the number of irreducible components of $H|_{S}$, then we have
@@
$\varphi\colon(x,y,z,s)\mapstochar\rightarrow(x^{\prime},y^{\prime},z^{\prime},s^{\prime})$
@@
$\psi\colon(x,y,z,s)\mapstochar\rightarrow(x,y,z,s^{\prime})$
@@
By Lemma 6.1, there is an open neighborhood $W$ of $\mathbf{0}_{3}$ in $\overline{W}$, there are open neighborhoods $T\subseteq T^{\prime}$ of $o$ in $\overline{T}$, such that the restriction of $s^{\prime}$ on $W\times T$ is deformation equivalent to the projection function $(x,y,z,s)\mapstochar\rightarrow s$, with values inside $T^{\prime}$. Up to shrinking $W$ and $T^{\prime}$, we may assume that $W\times T^{\prime}\subseteq Z$. As a consequence, the image $\psi(W_{sm}\times\{o\})$ is deformation equivalent to $W_{\operatorname{\textsubscript{\rm sm}}}\times\{o\}$ inside $Z_{\operatorname{\textsubscript{\rm sm}}}=Z\setminus(\{\mathbf{0}_{3}\}\times S^{\circ})$. Then by Lemma 6.2, the image of
$\psi_{*}\colon\pi_{1}(W_{\operatorname{\textsubscript{\rm sm}}}\times\{o\})\rightarrow\pi_{1}(Z_{\operatorname{\textsubscript{\rm sm}}})$